The Costaguanan stock market provided a rate of return of 98%. The inflation rate in Costaguana during the year was 86%. In the Ruritania, in contrast, the stock market return was only 15%, but the inflation rate was only 3%. Calculate the real rate of return for both the Costaguana and the Ruritania stock markets.

a)

Rate of Return
Costaguana
Ruritania

b) Which country’s stock market provided the higher real rate of return?
Answer #1

1.
Costaguana=1.98/1.86-1=6.4516%

Ruritania=1.15/1.03-1=11.6505%

2.
Ruritania's stock market provided the higher real rate of return